# villa-X: Enhancing Latent Action Modeling in Vision-Language-Action Models

### v2

@@ 4.1 Does our improved LAM learn higher-quality latent actions?
We are interested in how closely the predicted actions match the ground truth. In LIBERO, the robot action space has 8 dimensions: 3 for position, 4 for rotation, and 1 for the gripper. Rather than averaging the error across dimensions, we focus on the maximum L1 error across all action dimensions, as we want to avoid large deviations in any single aspect of the action. For each model variant (w/pp and wo/pp), we compute the number of validation samples whose maximum L1 error falls below a threshold. By sweeping this threshold, we count how many samples fall within each error bin. A better model should yield more samples with low errors.
@@
For each error bin, we compute the difference in the number of samples between the w/pp and wo/pp variants and present the results in Figure 4. The w/pp variant produces more samples with smaller errors, while the wo/pp variant has more samples in the high-error bins. This demonstrates the effectiveness of the proprio FDM module in capturing information from the robot actions.
@@
Table 1: Evaluation results on SIMPLER for different variants of our villa-X (top group) and alternative approaches for incorporating latent actions (bottom group). ”Ours” refers to the w/pp described in the main text.
@@
Method Google robot WidowX robot Pick Move Drawer Place Avg. Carrot Eggplant Spoon Cube Avg. Ours 81.7 55.4 38.4 4.2 44.9 24.2 71.7 48.3 19.2 40.8 wo/pp 77.0 52.7 42.6 2.8 43.8 22.5 57.5 43.3 5.9 32.3 wo/LAM 42.1 24.6 38.4 0.0 26.3 25.8 60.8 36.7 9.2 33.1 LAPA-style 64.7 28.8 38.0 5.6 34.2 0.8 0.0 2.5 0.8 1.0 Go-1-style 29.0 38.0 31.3 4.6 25.7 5.8 50.8 1.7 1.0 14.8
@@
Policy Pre-training  Next, we compare how the latent actions generated by different variants of LAM (w/pp and wo/pp) influence policy pre-training. Unlike the main experiments, we pretrain models in this section on a mixture of 10% Fractal [6] data, 10% Bridge V2 [18] data, and 100% Something-Something V2 [20] data, to reduce computation cost while remaining a setting where limited robot data are available for training the VLA model. The resulting policies are evaluated in the SIMPLER environment [32], a simulation benchmark explicitly designed to mitigate the gap between simulated and real‐world robotic environments. It comprises two platforms, namely Google robot and WidowX robot, each with four manipulation tasks. We evaluate our method on all the 8 tasks on the visual matching setting. The results are summarized in Table 1. We observe that w/pp clearly outperforms wo/pp, demonstrating the effectiveness of incorporating the proprio FDM module. Additionally, we include a baseline that does not use latent actions (denoted wo/LAM) and is trained solely to predict robot actions. The performance of wo/LAM is significantly worse, indicating that pre-training with latent actions is essential.
LAM visualization
Figure 2 visualizes image pairs sharing the same latent action, demonstrating that these pairs correspond to similar underlying robot behaviors. To further demonstrate the transfer ability of our LAM, we extract latent actions from arbitrary video sequences, map them to robot actions using the proprio FDM, and execute the resulting robot action in SIMPLER simulator. The simulated motions closely reproduce the original demonstrations, indicating that latent actions are both aligned with and grounded in the robot’s actions. For more examples, please check Appendix C.1.
Figure 5: Transfer of human demonstrations to new scenes via latent actions and a world model. (a) moving right. (b) moving downwards. Please visit our website for more demos.
Figure 6: Transfer of video demonstrations into robot action executions via latent actions and a proprio FDM. Please visit our website for more demos.
LAM transfer consistency
We provide two visualizations to further evaluate the transfer consistency of the learned latent actions. Specifically, we first use the trained LAM to extract a sequence of latent actions from a source video. These latent actions are then applied to different initial images, allowing a world model to generate subsequent frames conditioned on these actions. As shown in Figure 5, the transfer rollouts (highlighted in blue) successfully identify the movable parts and demonstrate behaviors consistent with the original video.
We also evaluate the alignment between latent and robot actions. To do this, we use the proprio FDM to decode latent actions into executable robot actions. These actions are then executed in the SIMPLER simulator, with the resulting rollouts visualized in Figure 6. In each example, the top row shows the original video demonstration from which latent actions were extracted, while the bottom row displays the corresponding simulated execution. The simulated motions closely match the original demonstrations, confirming that the latent actions learned by villa-X are well-aligned with and grounded in the robot actions.
4.2 Can ACT-latent successfully plan future motions?
Figure 7: Transfer of video demonstrations into robot action executions via latent actions and a proprio FDM. Please visit our website for more demos.
In this experiment, we demostrate the motion planning capabilities of ACT-latent by visualizing its planned actions with a world model. For a given initial image and language instruction, ACT-latent generates a sequence of latent actions, and the world model takes the inital frame and latent actions as inputs to render the latent actions into planned future videos. Figure 7 shows results for both in-distribution and out-of-distribution samples, where the in-distribution samples are randomly sampled from the validation set but from the same datasets as used in training, and the out-of-distribution samples are from new real-world scenarios. The results demonstrate that ACT-latent successfully follows the language instructions to solve the tasks, while accurately identifying target objects and generating latent actions that successfully follow the instructions. ACT-latent also successfully identifies the concepts in the emoji, which rarely appears in robot datasets, suggesting that villa-X keeps the general vision-language capabilities in the initial VLM model after pre-training.
@@ 4.3 Can the actor module effectively leverage the pre-trained latent actions?
Following the experiment setup in the previous subsection, we train all models on the same dataset mixture and then evaluate the resulting policies in the SIMPLER environment [32]. The results are shown in Table 1. Compared to other two approaches, our method achieves significantly higher performance, validating the effectiveness of our design for incorporating latent actions into VLA pre-training.
@@ 4.4.1 SIMPLER Benchmark
Method Google Robot WidowX Robot Pick Move Drawer Place Avg. Carrot Eggplant Spoon Cube Avg.. RT-1-X * 56.7 31.7 59.7 21.3 42.4 4.2 0.0 0.0 0.0 1.1 Octo-base * 17.0 4.2 22.7 0.0 11.0 8.3 43.1 12.5 0.0 16.0 OpenVLA * 16.3 46.2 35.6 0.0 24.5 0.0 4.1 0.0 0.0 1.0 RoboVLMs * 72.7 66.3 26.8 36.1 50.5 25.0 0.0 20.8 8.3 13.5 RoboVLMs 77.3 61.7 43.5 24.1 51.7 20.8 79.2 45.8 4.2 37.5 GR00T-N1.5 69.3 68.7 35.8 4.0 44.5 54.3 61.3 75.3 57.0 62.0 MoTo 74.0 60.4 43.1 N/A N/A N/A N/A N/A N/A N/A LAPA N/A N/A N/A N/A N/A 45.8 58.3 70.8 54.2 57.3 Ours w/o latent 56.3 25.8 27.3 13.9 30.8 31.3 74.6 61.7 28.3 49.0 Ours 98.7 75.0 59.3 5.6 59.6 46.3 64.6 77.9 61.3 62.5
@@
We use the SIMPLER benchmark as described above. In this section, we compare against two categories of prior work:
  • Vision-Language-Action (VLA) models: RT-1-X [13], Octo-base [54], OpenVLA [30], RoboVLMs [34], which learn policies solely from mixed robot datasets.
@@
Except where noted (*), all models follow a two‐stage pretraining–finetuning protocol, including a general pretraining phase on large‐scale mixed data, followed by finetuning on a dataset of specific embodiment. We also include an ablation (villa-X w/o latent) that removes our latent‐action expert while keeping all other components unchanged. Reported baseline scores are taken directly from the original publications, while missing entries are marked as N/A.
@@
Table 2 summarizes the success rates on both platforms. Our full model achieves the highest score on average success rate on the Google robot (59.659.6%) and the WidowX robot (62.562.5%). This improvement over VLA methods, which cannot exploit unlabelled video, demonstrates the benefit of our incorporating human videos into policy learning. Moreover, our approach outperforms other latent‐action methods, indicating that our specific mechanism for leveraging video data is more effective. Finally, the gap between our full model and the “villa-X w/o latent” ablation confirms that the latent‐action expert is essential for achieving these gains.
4.4.2 LIBERO Benchmark
Baselines and Experimental Setup
We compare with the following existing models: Diffusion Policy [12] trained from scratch, Octo [54], and OpenVLA [30]. All models follow a two-stage pretraining-finetuning protocol. We finetune villa-X and villa-X w/o latent on the demonstration data of the each task suite separately, and test on the LIBERO simulator for 10 tasks and 20 trials per task on each task suite.
Experimental Results
Table 3 summarizes the success rates on each task suite of LIBERO. Our model achieves better performance than existing methods in all the four task suites. Also, our model with latent action achieves higher performance on 3 of the 4 task suites and average performance, confirming that the proposed latent action expert improves the manipulation performance.
Method Spatial Object Goal Long Average Diffusion Policy [12] 78.3 92.5 68.3 50.5 72.4 Octo-base [54] 78.9 85.7 84.6 51.1 75.1 OpenVLA [30] 84.7 88.4 79.2 53.7 76.5 Ours w/o latent 86.0 86.5 85.0 70.0 81.9 Ours 97.5 97.0 91.5 74.5 90.1
@@ 4.5 Evaluating villa-X on Real-world Robots
Figure 8: Real-world robot evaluation platforms: (top) Realman robot arm platform with a gripper and (bottom) Xarm robot arm with Xhand dexterous hand. Platform setups are shown on the left, with corresponding evaluation tasks on the right. Please visit our website for more videos.
@@
To further assess generalization, we evaluate our approach on two real‑world robotic platforms: a Realman robot arm with a gripper, and an Xarm manipulator equipped with a 12‑DOF Xhand dexterous hand. See Figure 8 for details.
Realman robot arm with gripper
On the robot-arm manipulation platform, we use a 6-DoF Realman RM 75 robot arm and a 1-DoF Inspire gripper. We fine-tune and evaluate our policy on “Pick-in” (pick the block into a bowl), “Pick-out” (pick the block out of a bowl), “Stack” (stack the block onto another block), “Unstack” (unstack the block from another block), and “Push” (push the block to a given location) tasks. For fine-tuning, we collect a dataset of 375 trajectories through teleoperation (with 75 trajectories for each task), where the object layout and table setup are fixed and only the object locations are dynamic.
We conduct two sets of evaluation: In task evaluation, we remain the table setup the same as data collection; in generalization evaluation, we change the color of the block and table cover. We evaluate the policy with 10 trials per task with each trial conducted with different object location, where all the experiment settings are the same for the evaluation of different policies (including the object location and lightening condition). The results are shown in Table 5, demonstrating that villa-X outperforms existing baselines in both task and generalization evaluation sets.
@@ 5 Conclusion, Limitations, and Future Works
In this paper, we presented villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that improves both the learning of latent actions and their incorporation into VLA pre-training. Our experiments demonstrate that our enhanced Latent Action Model learns higher-quality latent actions, and our improved policy model more effectively leverages these learned actions. Overall, our method exhibits superior performance in both simulated environments and real-world robotic tasks.
@@
One limitation is that the proposed latent expert, although effective at future planning through both visual and proprioceptive state planning, is not fully explored in this work. For example, future research could learn a critic with prior knowledge from foundation vision-language models, allowing multiple samples from the latent expert and rejecting planned trajectories that do not follow the language instruction. We leave this aspect as future work to further improve the capability of the ViLLA framework.
Appendix A Model Design
A.1 Latent Action Model
There are four modules in our latent action model. A ST-Transformer based inverse dynamic model (IDM), a vector quantization module, an image reconstruction forward dynamic model (FDM) and a robot proprio FDM. The IDM takes an 8 × 224 × 224 video clip as input, following a Patch Embedding with a patch size of 14 and 12 ST-blocks, where each block has a hidden dimension of 768 and 32 attention heads. The codebook size of vector quantization is set to 32. Image reconstruction FDM is a 12-layer ViT-base network, and robot proprio FDM is a 2-layer MLP with a hidden size of 768. Our model is trained with a learning rate of 1.5e-4, a batch size of 512, and a 2000-step linear warmup. We use the same loss weight for the image FDM and proprio FDM. The pretraining takes 4 days on 128 NVIDIA A100 GPUs.
A.2 Actor Module

### v3

@@ 4 Experiments
  • By scaling pre-training, can the latent actor module effectively plans for the future and generalize zero-shot to unseen embodiments and open-vocabulary concepts in symbolic icons?
@@ 4.1 Does our improved LAM learn higher-quality latent actions?
We are interested in how closely the predicted actions match the ground truth. In LIBERO, the robot action space has eight dimensions: three for position, four for rotation, and one for the gripper. Rather than averaging the error across dimensions, we focus on the maximum L1 error across all action dimensions, as we want to avoid large deviations in any single aspect of the action. For each model variant (w/pp and wo/pp), we compute the number of validation samples whose maximum L1 error falls below a threshold. By sweeping this threshold, we count how many samples fall within each error bin. A better model should yield more samples with low errors.
@@
For each error bin, we compute the difference in the number of samples between the w/pp and wo/pp variants and present the results in Figure 3. The w/pp variant produces more samples with smaller errors, while the wo/pp variant has more samples in the high-error bins. This demonstrates the effectiveness of the proprio FDM module in capturing information from the robot actions. We further visualize the learned latent actions, and perform more ablations on LAM. Please refer to Appendix D for more details.
@@
Table 1: Evaluation results on SIMPLER for different variants of our villa-X(top group) and alternative approaches for incorporating latent actions (bottom group). “Ours” refers to the w/pp described in the main text.
@@
Method Google robot WidowX robot Pick Move Drawer Avg. Carrot Eggplant Spoon Cube Avg. Ours 81.7 55.4 38.4 58.5 24.2 71.7 48.3 19.2 40.8 wo/pp 77.0 52.7 42.6 57.4 22.5 57.5 43.3 5.9 32.3 wo/LAM 42.1 24.6 38.4 35.0 25.8 60.8 36.7 9.2 33.1 LAPA-style 64.7 28.8 38.0 43.8 0.8 0.0 2.5 0.8 1.0 Go-1-style 29.0 38.0 31.3 32.8 5.8 50.8 1.7 1.0 14.8
@@
Policy Pre-training  Next, we compare how the latent actions generated by different variants of LAM (w/pp and wo/pp) influence policy pre-training. Unlike the main experiments, we pretrain models in this section on a mixture of 10% Fractal [5] data, 10% Bridge V2 [17] data, and 100% Something-Something V2 [19] data, to reduce computation cost while remaining a setting where limited robot data are available for training the VLA model. The resulting policies are evaluated in the SIMPLER environment [32], a simulation benchmark explicitly designed to mitigate the gap between simulated and real‐world robotic environments. It comprises two platforms: the Google robot with three manipulation tasks and the WidowX robot with four. We evaluate our method on the visual matching setting. The results are summarized in Table 1. We observe that w/pp clearly outperforms wo/pp, demonstrating the effectiveness of incorporating the proprio FDM module. Additionally, we include a baseline that does not use latent actions (denoted wo/LAM) and is trained solely to predict robot actions. The performance of wo/LAM is significantly worse, indicating that pre-training with latent actions is essential.
@@ 4.2 Can the actor module effectively leverage the pre-trained latent actions?
Following the experiment setup in the previous subsection, we train all models on the same dataset mixture and then evaluate the resulting policies in the SIMPLER environment [32]. The results are shown in Table 1. Compared to other two approaches, our method achieves significantly higher performance, validating the effectiveness of our design for incorporating latent actions into VLA pre-training. More ablation studies on policy designs can be found in Appendix F.
4.3 Zero-shot Generalization for Latent Actor Module
Figure 4: Visualization of zero-shot latent plans on an unseen embodiment. Each pair of images shows the starting frame (left) and the ending frame (right), with the instruction displayed above.
To evaluate the zero-shot generalization capabilities of ACT-latent in planning, we conducted a real-world visualization experiment focused on its ability to handle new embodiments and understand novel open-vocabulary symbols. For embodiment generalization, we used a Realman robot arm, a new embodiment never seen during training. To assess open-vocabulary generalization, we designed a set of symbol cards, testing the model’s ability to comprehend concepts typically absent from standard robotics datasets.
The evaluation process is as follows: given a starting image and a language instruction (e.g., “touch the corn”), ACT-latent first generates a sequence of latent actions. Then a separately trained world model is used to render this sequence into a video, allowing us to verify the effectiveness of the plan.
As shown in Figure 4, the rendered trajectories confirm that the model successfully generates latent plans that follow the commands. These results highlight two key capabilities of our approach:
  • Embodiment Generalization: ACT-latent successfully identifies and controls this unseen robot arm, indicating that its learned knowledge is embodiment-agnostic and readily transferable to the new robotic platform.
  • Open-Vocabulary Understanding: The model’s ability to interact correctly with symbol concepts reveals that villa-X retains the general-purpose vision-language capabilities of the original Vision-Language Model (VLM) after pre-training.
More visualizations can be found in Appendix E. To evaluate how effectively villa-X uses this knowledge, we measure its success rates on a variety of control tasks in the following sections.
@@ 4.4 Evaluating villa-X in Simulation
Method Google Robot WidowX Robot Pick Move Drawer Avg. Carrot Eggplant Spoon Cube Avg.. RT-1-X * 56.7 31.7 59.7 49.4 4.2 0.0 0.0 0.0 1.1 Octo-base * 17.0 4.2 22.7 14.6 8.3 43.1 12.5 0.0 16.0 OpenVLA * 16.3 46.2 35.6 32.7 0.0 4.1 0.0 0.0 1.0 RoboVLMs * 72.7 66.3 26.8 55.3 25.0 0.0 20.8 8.3 13.5 RoboVLMs 77.3 61.7 43.5 60.8 20.8 79.2 45.8 4.2 37.5 $\pi_{0}$ 72.7 65.3 38.3 58.7 0.0 62.5 29.1 16.6 27.1 $\pi_{0}$-FAST 75.3 67.5 42.9 61.9 21.9 66.6 29.1 10.8 32.1 OpenVLA-OFT 72.3 69.6 47.2 63.0 4.2 N/A 12.5 8.3 N/A GR00T-N1.5 69.3 68.7 35.8 57.9 54.3 61.3 75.3 57.0 62.0 TraceVLA 45.0 63.8 63.1 57.3 16.6 65.0 12.5 16.6 27.7 Magma 75.0 53.0 58.9 62.3 29.2 91.7 37.5 20.8 44.8 MoTo 74.0 60.4 43.1 59.2 N/A N/A N/A N/A N/A LAPA N/A N/A N/A N/A 45.8 58.3 70.8 54.2 57.3 Ours w/o latent 56.3 25.8 27.3 36.5 31.3 74.6 61.7 28.3 49.0 Ours 98.7 75.0 59.3 77.7 46.3 64.6 77.9 61.3 62.5
@@
We use the SIMPLER benchmark as described above. In this section, we compare against several categories of prior work:
  • Vision-Language-Action (VLA) models: RT-1-X [12], Octo-base [52], OpenVLA [29], RoboVLMs [34], $\pi_{0}$ [4], $\pi_{0}$-FAST [54], OpenVLA-OFT [30], which learn policies solely from mixed robot datasets.
@@
  • Visual trace methods: TraceVLA [71], Magma [67], which learn planning on the extract visual traces from videos.
@@
Except where noted (*), all models follow a two‐stage pretraining–finetuning protocol, including a general pretraining phase on large‐scale mixed data, followed by finetuning on a dataset of specific embodiment. We also include an ablation (Ours w/o latent) that removes our latent‐action expert while keeping all other components unchanged. Baseline scores are cited from their original publications or other relevant literature, while missing entries are marked as N/A.
@@
Table 2 summarizes the success rates on both platforms. Our full model achieves the highest score on average success rate on the Google robot (77.777.7%) and the WidowX robot (62.562.5%). This improvement over VLA methods, which cannot exploit unlabelled video, demonstrates the benefit of incorporating human videos into policy learning. Moreover, our approach outperforms other video learning and latent‐action methods, indicating that our specific mechanism for leveraging video data is more effective. Finally, the gap between our full model and the “villa-X w/o latent” ablation confirms that the latent‐action expert is essential for achieving these gains.
@@ 4.5 Evaluating villa-X on Real-world Robots
To assess real-world generalization, we deploy villa-X on two platforms: a Realman arm with a gripper and an XArm with a 12-DoF XHand as shown in Figure 5.
Realman robot arm with gripper
We use a 6-DoF Realman RM 75 with a 1-DoF Inspire gripper, fine-tuning and evaluating on five tasks: Pick-in (pick the block into a bowl), Pick-out (pick the block out of a bowl), Stack (stack the block onto another block), Unstack (unstack the block from another block), and Push (push the block to a given location). The fine-tuning set contains 375 teleoperated trajectories (75 per task); the object layout and table are fixed, while object positions vary.
We conduct two sets of evaluation: In task evaluation, we remain the table setup the same as data collection; in generalization evaluation, we change the color of the block and table cover. For each task, we conduct 10 trials with distinct object positions; positions and lighting are identical across policies. As shown in Table 4, villa-X outperforms all baselines in both settings.
Figure 5: Real-world robot evaluation platforms: (top) Realman robot arm platform with a gripper and (bottom) Xarm robot arm with Xhand dexterous hand. Platform setups are shown on the left, with corresponding evaluation tasks on the right.
@@ 5 Conclusion, Limitations, and Future Works
In this paper, we presented villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that improves both the learning of latent actions and their incorporation into VLA pre-training. Our experiments demonstrate that our enhanced Latent Action Model learns higher-quality latent actions, and our improved policy model more effectively leverages these learned latents. The learned latent action expert can even generalize zero-shot to an unseen embodiment, showing strong generalization ability. Overall, our method exhibits superior performance in both simulated environments and real-world robotic tasks. One limitation is that the proposed latent expert, although effective at future planning through both visual and proprioceptive state planning, is not fully explored in this work. For example, future research could learn a critic with prior knowledge from foundation vision-language models, allowing multiple samples from the latent expert and rejecting planned trajectories that do not follow the language instruction. We leave this aspect as future work to further improve the capability of the ViLLA framework.
Appendix A Additional Implementation Details for LAM
In this appendix, we provide extended information on the architecture, training protocols, and inference behavior for our Latent Action Model (LAM).
A.1 Architecture Overview
Our LAM comprises four main modules:
  1. (i)
    Spatial–Temporal Transformer (ST-Transformer) Inverse Dynamics Model (IDM): Takes a video clip (by default, $8\times 224\times 224$) as input. We employ patch embedding with a patch size of $14$ and stack $12$ ST-blocks [65], each with a hidden dimension of $768$ and $32$ attention heads.
  2. (ii)
    Vector Quantization (VQ) Module: Maps the continuous IDM outputs to discrete latent tokens, each associated with a codebook entry. We set the codebook size to $32$. While the model internally uses discrete token indices during training, the continuous codebook centers are used in downstream modules.
  3. (iii)
    Image Reconstruction Forward Dynamics Model (FDM): A $12$-layer Vision Transformer (ViT)-base network that takes the current frame $o_{t}$ and a latent action $z_{t}$ to predict $\hat{o}_{t+K}$.
  4. (iv)
    Proprioceptive Forward Dynamics Model (proprio-FDM): A 2-layer MLP with dual output heads to predict future robot states $\hat{q}_{t+i}$ and low-level robot actions $\hat{a}_{t+i}$. This module takes the current robot state $q_{t}$, the latent $z_{t}$, and an embodiment context vector $\mathbf{c}_{e}$.
Rather than predicting a single latent action per pair of frames, the ST-Transformer-based IDM processes a sequence of $T_{\text{LAM}}$ frames, resulting in $T_{\text{LAM}}-1$ latent tokens. We use $T_{\text{LAM}}=8$. By reconstructing future frames with the FDM and future states/actions with the proprio-FDM, the model learns a latent representation that is both visually and physically grounded.
A.2 Training Details
We train our LAM on a combination of human egocentric data (e.g., Ego4D [21]) and robot trajectories (e.g., OpenX [12]). Samples lacking low-level robot annotations (e.g., human videos) exclude the proprio-FDM branch, using only the visual FDM objective.
Hyperparameters.
We use a batch size of $512$ and a learning rate of $1.5\times 10^{-4}$, with a $2000$-step linear warmup. Training lasts approximately $4$ days on $128$ NVIDIA A100 GPUs. Both the visual FDM and proprio-FDM share the same weighting in the overall loss. Throughout training, each latent token is discretized via the VQ module but is represented by its continuous codebook center in subsequent network components.
A.3 Inference Behavior and Diagnostics
During inference, only the IDM is required to extract latent tokens from consecutive frames. The FDM and proprio-FDM are typically retained for diagnostic and visualization purposes, allowing us to examine whether the learned latent tokens accurately capture future frame content, robot states, and actions. This reconstruction-based analysis aids in understanding and debugging the physical grounding of the latent representation.
Appendix B Additional Implementation Details for Actor Module